OpenStudy (anonymous):

PLEASE HELPP MEEE! Find the x-intercepts of the graph of each function: y=x^2+2x

OpenStudy (accessdenied):

Do you understand what the term "x-intercept" means, like in terms of the graph?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Do i factor the x^2 + 2x?

OpenStudy (accessdenied):

Once we have the equation (y=0 for x-intercepts) x^2 + 2x = 0 Yes, we can factor here to solve for the x-values.

OpenStudy (anonymous):

ok let me work it out

OpenStudy (anonymous):

x(x+2)

OpenStudy (accessdenied):

Yep, that's correct. Now, we just equate each factor to 0 to solve (If ab = 0,. then a= 0 or b= 0)

OpenStudy (anonymous):

y = x^2 + 2x y = x(x+2) x = 0 or x+2 = 0 x = -2

OpenStudy (anonymous):

i got x=0 or x=-2!

OpenStudy (anonymous):

if finding x intercepts , then you make it so that y = 0 and then solve (that's what i did) y = x^2 + 2x 0 = x^2 + 2x 0 = x(x+2) and so then you get x = 0 and -2

OpenStudy (anonymous):

thank you both so much this really helped a lot!

OpenStudy (anonymous):

if you ever have to find y -intercepts, you set X equal to 0 as opposed to Y. no problem (:

OpenStudy (accessdenied):

You're welcome. :)

OpenStudy (anonymous):

Thank youu!